# Energy barriers for small electron polaron hopping in bismuth ferrite from first principles

### v1

@@ Abstract
Evidence from first-principles calculations indicates that excess electrons in BiFeO3 form small polarons with energy levels deep inside the electronic band gap. Hence, -type electronic transport could occur by hopping of small electron polarons rather than by band-like transport. Here, by means of first-principles calculations, small electron polaron hopping in BiFeO3 is investigated. Both bulk BiFeO3 and a typical ferroelectric domain wall, the neutral 71° domain wall, are considered. The latter is included to account for experimental observations of electrical conductivity at domain walls in otherwise insulating ferroelectrics. The object of this study is to shed light on the intrinsic electron conduction in rhombohedral BiFeO3 and the effect of pristine neutral ferroelectric domain walls. The computed energy barriers for small electron polaron hopping are of the order of the thermal energy at room temperature, both in bulk and within the neutral 71° domain wall. The domain wall is found to act as a two-dimensional trap for small electron polarons with a trap depth of about two to three times the thermal energy at room temperature. Based on these findings, the intrinsic -type mobility and diffusion constant in BiFeO3 at room temperature are estimated, and experimental conductivity data for BiFeO3 are discussed.
@@ 1 Introduction
Like hematite, BiFeO3 might be suitable for water splitting. Other than hematite, BiFeO3 is ferroelectric and can form ferroelectric domain walls, where electronic properties differ from those inside the bulk. In particular, electrical conductivity has been observed at ferroelectric domain walls, whereas the surrounding bulk material is insulating or semiconducting [9, 10, 11, 12, 13, 14]. Like in hematite [15, 16], there is computational evidence that excess electrons in BiFeO3 form small polarons with energy levels deep inside the band gap [17, 18], and that the small electron polarons are trapped by ferroelectric domain walls [19, 20, 21]. Hence, electronic transport in -doped BiFeO3 could occur by hopping of small electron polarons rather than by band-like transport, and the hopping of electron polarons could require less energy than that needed to promote an electron from the polaron state to the conduction band, see Figure 1(b). Domain walls, if present, could affect electronic transport. Here, first-principles electronic structure calculations based on density-functional theory were performed to determine the energy barriers for electron hopping in the bulk and at one of the prevalent ferroelectric domain walls, the neutral 71° domain wall [22, 23]. The 71° domain wall was chosen because it affects the atomic and electronic structure more strongly than the 109° wall [22, 20], and because it has been reported to be conductive in several studies, such as Ref. [13, 24, 25]. The object of this study is to shed light on the intrinsic -type conduction mechanism in BiFeO3 and on the effect of pristine neutral ferroelectric domain walls. The approach taken here allows to determine the intrinsic mechanism of -type conduction, once excess electrons have been brought into BiFeO3 in some way. For example, charge carriers may originate from point defects, such as oxygen vacancies or aliovalent dopants or impurities, from photoexcitation resulting in bound or dissociated electron-hole pairs, or from an electrical contact. The approach taken here can predict the intrinsic charge carrier mobility, but it cannot predict the charge-carrier density since the latter depends on the sample fabrication and history and the experimental conditions, and it can hence not predict the absolute electrical conductivity, but it can predict the temperature dependence of the electrical conductivity and the effect of deviations from the perfect crystal structure. External processes of emitting and trapping electrons by defects are not part of the intrinsic electrical conduction mechanism, and considering the effects of all possible defects would be beyond the scope of this work. However, a particular defect (or desired structure modification) is considered here, namely the neutral ferroelectric 71° domain wall.
This paper is organized as follows: First the computational methods are introduced, then the energy barriers for small polaron hopping are presented for bulk and domain wall and the electronic mobility and diffusion constant at room temperature are determined. Finally results are discussed based on experimental data, where available.
@@ 2 Methods
The calculations were performed with the program vasp [26], using the Projector-Augmented Wave (PAW) method [27, 28] and pseudopotentials with 5 (Bi), 16 (Fe), and 6 (O) valence electrons, respectively. Periodic boundary conditions were employed. The local spin-density approximation (LSDA) to density-functional theory (DFT) was used, and the band gap was corrected with a Hubbard- of 5.3 eV applied to the Fe- states using Dudarev’s scheme [29]. This approach yields atomic and electronic structure in close agreement with experiment [19, 20, 21]. Standard approaches to model polarons and polaron hopping from first principles are DFT+ or hybrid DFT functionals, which were applied, e.g., in Ref. [30, 31]. Here, to test how much the calculated hopping barriers depend on the level of theory, the electronic structure with polaron and the polaron hopping barriers in bulk BiFeO3 obtained with LSDA+ and with the hybrid HSE functional [32] for the geometries optimized with LSDA+ are compared. A similar comparison of the hybrid PBE0 functional [33] and DFT+ and a piece-wise linear DFT functional was made in Ref. [34] for polarons in Ga2O3, with the result that polaron formation energy and eigenvalue from DFT+ and PBE0 agree at least semiquantitatively. Plane-wave basis functions with energies up to 520 eV were used. Both the atomic positions and the cell parameters were optimized until the total energy differences between consecutive iteration steps fell below 0.01 meV for the optimization of the electronic density and below 0.1 meV for the optimization of the atomic structure. BiFeO3 with excess electrons was modeled using supercells that comprise eight (bulk) or 16 (domain wall) 10-atom unit cells, respectively. Convergence of energy barriers with respect to supercell size was confirmed for bulk, see Supporting Information. One excess electron was added to the supercell, whose charge was compensated by a uniform background charge density. The magnetic structure of BiFeO3 was approximated by that of a collinear -type antiferromagnet, spin-orbit coupling was neglected.
@@ 2.1 Excess electrons in bulk
BiFeO3 bulk with excess electrons was modeled using supercells with 80 atoms that consist of rhombohedral 10-atom unit cells. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length, as depicted in Figure 2(a).
2.2 Excess electrons at the ferroelectric 71° domain wall
Excess electrons at 71° domain walls were modeled using a 160-atom supercell, see Figure 2(b), spanned by 4() ( in the pseudocubic system, here “”), (- ( in the pseudocubic system), and () ( in the pseudocubic system), where , , and span the pseudocubic 5-atom cell. This supercell is based on that used in Ref. [19, 20, 21]. The atomic structures and formation energies of low-energy ferroelectric domain walls in BiFeO3 are well known [35, 22, 36, 23, 37]. Atomic and electronic structure of the neutral 71° domain wall investigated here have been published elsewhere [23, 22, 19, 20, 21]. The antiferromagnetic -type spin configuration of the bulk was maintained in the systems with domain walls. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length.
@@ 2.3 Polaron hopping
Energy barriers for polaron hopping between different Fe sites were obtained using the nudged-elastic-band (NEB) method, in which an artificial spring force acting along the hopping path constrains atomic structure optimization to directions perpendicular to the path. Nine configurations (NEB images) along each hopping path were considered, including initial and final state.
Both hops between nearest-neighbor (NN) and second-nearest neighbor (2NN) Fe sites were considered, since NN hops require spin flip of the electron and should hence be suppressed.
The number of inequivalent NN and 2NN hops is limited by symmetry. In bulk, there is one NN hop and two inequivalent 2NN hops: a 2NN hop perpendicular to the ferroelectric polarization direction, and a 2NN hop in a direction that forms an angle of ° with the polarization direction, see Figure 2(a). At ferroelectric domain walls, Fe sites are still equivalent within planes parallel to the wall, but differ between these planes. Hops can occur along the wall or with a component perpendicular to the wall, see Figure 2(b). NEB calculations were considered converged when the energy barrier changed by maximally 1 meV in two consecutive iterations of the NEB structure optimization. The strain dependence of the barriers was also investigated. The strain reference (zero strain, ) corresponds to the optimized geometry of bulk or domain wall without excess electron. Tensile strain, , corresponds to the optimized geometry of bulk or domain wall with excess electron.
@@ 2.4 Transition state theory
is the activation energy (hopping barrier) for the 2NN hop, is the reorganization energy. For symmetric hops, . The electronic coupling matrix elements , where the are electronic wave functions localized on the initial and final Fe site, respectively, were adopted from Ref. [40], where they were calculated for Fe(H2O) complexes as a function of the Fe-Fe distance from a model Hamiltonian that treated Fe with a pseudopotential and H2O as classical point charges. Here, the matrix element for the second-nearest Fe-Fe distance calculated for bulk BiFeO3, 5.516 , was used throughout (-101.875 cm eV). Alternatively, the coupling matrix elements could be determined as the deviation of the calculated barrier from the parabolic crossing point, see Figure 3(a). In the case of compressive strain, where the calculated hopping energy profile is close to a double parabola, one obtains a coupling matrix element of to meV.
@@
where is the elementary charge, is the hopping distance, is the number of equivalent available final states for a given initial state, is the dimension, is Boltzmann’s constant, and is the temperature, here 300 K. For bulk, =3, =12 for 2NN hops, is the 2NN Fe-Fe distance, and the pseudocubic lattice constant (about 4 Å). For two-dimensional transport within a domain-wall plane, . Hops along the coordinate consist of two consecutive hops, such as , with and . Hops along the coordinate can be single 2NN hops with and =2, such as , or double 2NN hops with and , such as . The effective barrier for double hops was taken as the higher one of the two. The two-dimensional -type mobility and the diffusion constant in the domain wall are computed as
@@ 3.1 Excess electrons in bulk
The electronic density of states (DOS) for a supercell (80 atoms) with an electron polaron is shown in Figure 3(b). The DOS obtained with LSDA+ and HSE agree semiquantitatively.
The computed polaron level lies 0.7 eV below the conduction band minimum for LSDA+ and 1 eV below the conduction band minimum for HSE,
with an estimated uncertainty of 0.2 eV.
[…]
The energy barriers for NN hops and the two different 2NN hops are depicted in Figure 4(a) and (b) for different strains . Both NN and 2NN barriers are of the order of the thermal energy at room temperature and shrink upon compression. Hence, the hopping barrier is an order of magnitude smaller than the energy distance to the conduction band minimum. The 2NN barriers are not strongly direction dependent. The transitions from initial to final site are non-adiabatic (abrupt), as seen from the isosurface of the excess electron’s density as a function of the hopping coordinate (yellow orbitals in Figure 4(c)–(f)). The computed energy barriers are listed in Table 1.
NN | 2NN, ° | 2NN, ° | ||||
LSDA+ | HSE | LSDA+ | HSE | LSDA+ | HSE | |
% | 8 | 6 | 17 | 18 | 16 | 14 |
0 | 15 | 18 | 24 | 34 | 20 | 20 |
0.64% | 15 | 25 | 27 | 39 | 29 | 30 |
To obtain the electron mobility and the diffusion constant at room temperature, the barrier at 0.65%, corresponding to the experimental lattice constant at room temperature [41], was used. With an energy barrier of 28 meV (the average of the two 2NN barriers) and a hopping distance of the 2NN Fe-Fe distance, the computed electron mobility in BiFeO3 bulk at room temperature is 0.6 cm2/Vs, the computed diffusion constant is 0.02 cm2/s. The uncertainty of the computed and are about one order of magnitude.

### v2

# Energy profile and hopping barriers for small electron polarons at ferroelectric domain walls in bismuth ferrite from first principles

@@ Abstract
Evidence from first-principles calculations indicates that excess electrons in BiFeO3 form small polarons with energy levels deep inside the electronic band gap. Hence, -type electronic transport could occur by hopping of small electron polarons rather than by band-like transport. Here, by means of first-principles calculations, small electron polaron hopping in BiFeO3 is investigated. Both bulk BiFeO3 and a typical ferroelectric domain wall, the neutral 71° domain wall, are considered. The latter is included to account for experimental observations of electrical conductivity at domain walls in otherwise insulating ferroelectrics. The object of this study is to shed light on the intrinsic electron conduction in rhombohedral BiFeO3 and the effect of pristine neutral ferroelectric domain walls. The computed energy barriers for small electron polaron hopping are near 0.2 eV, similar to other perovskite oxides, both in the bulk and within the neutral 71° domain wall. Trapping energies of small electron polarons at the three prevalent domain walls, the 71°, the 109°, and the 180° wall, were determined. The domain walls are found to act as two-dimensional traps for small electron polarons, with a trap depth of about two times the thermal energy at room temperature. Based on these findings, the intrinsic -type mobility and the diffusion constant in BiFeO3 at room temperature are estimated, and experimental conductivity data for BiFeO3 are discussed.
@@ 1 Introduction
Like hematite, BiFeO3 might be suitable for water splitting. Other than hematite, BiFeO3 can form ferroelectric domain walls, where electronic properties differ from those inside the bulk. In particular, electrical conductivity has been observed at ferroelectric domain walls in BiFeO3 and in other ferroelectric perovskite oxides, whereas the surrounding bulk material is insulating or semiconducting [9, 10, 11, 12, 13, 14]. Like in the case of hematite [15, 16], there is computational evidence that excess electrons in BiFeO3 form small polarons with energy levels deep inside the band gap [17, 18], and that the small electron polarons are trapped by ferroelectric domain walls [19, 20, 21]. Hence, electronic transport in -doped BiFeO3 should occur by hopping of small electron polarons rather than by band-like transport, if hopping of electron polarons requires less energy than promoting an electron from the polaron state to the conduction band, see Figure 1(b). Ferroelectric domain walls, if present, could affect electronic transport.
Here, for the first time, first-principles electronic structure calculations based on density-functional theory were performed to determine the energy barriers for electron polaron hopping in BiFeO3. Different hopping directions were considered so that the anisotropy of the crystal structure was fully taken into account. For the first time, polaron hopping at a ferroelectric domain wall was modeled from first principles. The neutral 71° domain wall was chosen because it is one of the prevalent ferroelectric domain walls in BiFeO3 [22, 23], because it affects the atomic and electronic structure more strongly than the 109° wall [20, 22], and because it has been reported to be conductive (about one order of magnitude more than the domain interior) in several studies, such as Ref. [13, 24, 25]. The object of this study is to shed light on the intrinsic -type conduction mechanism in BiFeO3 and on the influence of pristine neutral ferroelectric domain walls.
This paper is organized as follows: First the computational methods are introduced, then the energy landscapes and hopping barriers for small polarons are presented for bulk and domain wall, and the electronic mobility and the diffusion constant at room temperature are determined. Finally results are discussed based on experimental data, where available.
@@ 2 Methods
The calculations were performed with the “Vienna Ab initio Simulation Package” (Vasp) [26], using the Projector-Augmented Wave (PAW) method [27, 28] and pseudopotentials with 5 (Bi), 16 (Fe), and 6 (O) valence electrons, respectively. Periodic boundary conditions were employed. The local spin-density approximation (LSDA) to density-functional theory (DFT) was used, and the band gap was corrected with a Hubbard- of 5.3 eV applied to the Fe- states using Dudarev’s scheme [29]. This approach yields atomic and electronic structure in close agreement with experiment [19, 20, 21], see also Table 1.
Standard approaches to model polarons and polaron hopping from first principles are DFT+ or hybrid DFT functionals, which were applied, e.g., in Ref. [30, 31]. Here, to test how much the calculated hopping barriers depend on the level of theory, the electronic structure with polaron and the polaron hopping barriers in bulk BiFeO3 from LSDA+ are compared to those from the hybrid HSE06 functional [32] (25% Hartree-Fock exchange, 5 Å screening length) for the geometries optimized with LSDA+. A similar comparison of the hybrid PBE0 functional [33] and DFT+ and a piece-wise linear DFT functional was made in Ref. [34] for polarons in Ga2O3, with the result that polaron formation energy and eigenvalue from DFT+ and PBE0 agree at least semiquantitatively.
Plane-wave basis functions with energies up to 520 eV were used. Both the atomic positions and the cell parameters were optimized until the total energy differences between consecutive iteration steps fell below 0.01 meV for the optimization of the electronic density and below 0.1 meV for the optimization of the atomic structure.
@@ 2.1 Excess electrons in bulk
BiFeO3 bulk with excess electrons was modeled using supercells with 80 atoms that consist of rhombohedral 10-atom unit cells. Convergence of energy barriers with respect to the supercell size was confirmed by comparing with results obtained with a 40-atom supercell, see Supp. Inf. The barriers obtained with the two supercell sizes differ only by about 10%. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length, as depicted in Figure 2(a). Convergence with respect to -point density was confirmed by comparing with results obtained with a mesh, see Supp. Inf. One excess electron was added to the supercell, whose charge was compensated by a uniform background charge density. The magnetic structure of BiFeO3 was approximated by that of a collinear -type antiferromagnet, spin-orbit coupling was neglected.
2.2 Excess electrons at the ferroelectric 71°, 109°, and 180° domain wall
The atomic structures and formation energies (here 172 mJ/m2 for the 71° wall, 63 mJ/m2 for the 109° wall, and 86 mJ/m2 for the 180° wall [20]) of low-energy ferroelectric domain walls in BiFeO3 are well known [22, 23, 35, 36, 37]. Atomic and electronic structure of the neutral 71°, 109°, and 180° domain walls investigated here have been published elsewhere [19, 20, 21, 22, 23]. The computed properties of the ferroelectric 71°, 109°, and 180° domain walls, such as domain-wall energies and widths, can be found in the Supplemental Material to [20]. The antiferromagnetic -type spin configuration of the bulk was maintained in the systems with domain walls. The energy profile of small electron polarons as a function of the distance from the domain walls was determined and the trapping energy at the domain walls was obtained. 240-atom supercells were employed with 12 perovskite layers in the direction perpendicular to the walls (six layers per domain, see Figure 6). -point meshes of -points were used. Very good convergence of results was obtained with this mesh, compared to a -point mesh, see Supporting Information. The energy convergence thresholds were set to eV for the electronic structure and to eV for the atomic coordinates, respectively, which is probably one order of magnitude tighter than necessary. The trapping energy was determined as the total energy difference between a polaron position in the domain interior (DI) and in the lowest-energy configuration at the domain wall (DW), . Energy profiles and trapping energies were corrected by a factor , see Table 1, to avoid underestimation of energy differences that originate in ferroelectricity. Based on the computed trapping energies, the concentration of polarons was estimated based on a Boltzmann distribution,
where is the concentration far away from the domain wall, is the total energy difference between a polaron in and in the domain interior, is Boltzmann’s constant, and is the temperature.
Hopping of small electron polarons at 71° domain walls was modeled using a 160-atom supercell, see Figure 2(b), spanned by 4() ( in the pseudocubic system, here “”), (- ( in the pseudocubic system), and () ( in the pseudocubic system), where , , and span the pseudocubic 5-atom cell. This supercell is based on that used in Ref. [19, 20, 21]. The first Brillouin zone of the supercell was sampled with -points, equivalent to -points for the primitive rhombohedral ten-atom unit cell of BiFeO3 and to -points for a pseudocubic five-atom perovskite cell with about four Å edge length.
@@ 2.3 Polaron hopping
Energy barriers for polaron hopping between different Fe sites were obtained using the nudged-elastic-band (NEB) method, in which an artificial spring force acting along the hopping path constrains atomic structure optimization to directions perpendicular to the hopping path. Nine configurations (NEB images) along each hopping path were considered, including initial and final state. The NEB method is designed to determine the minimum-energy path, which can have a considerably lower energy barrier than the shortest path (the linear path). Since here the top of the barrier is pointed (the forces along the minimum energy path are not continuous), which might pose numerical difficulties for the NEB algorithm, the top of the barrier was obtained by calculating the energy along a linear path between the two nearest NEB images (images 4 and 5). The deviation of the barrier from the highest NEB energies obtained for the 2NN hops in bulk were then used to estimate the barrier tops for hops at the domain wall, i.e. the energy of the linear path (between images 4 and 5) was not calculated for the domain-wall hops. Both hops between nearest-neighbor (NN) and second-nearest neighbor (2NN) Fe sites were considered, but NN hops require spin flip of the electron and should hence be suppressed.
The number of inequivalent NN and 2NN hops is limited by symmetry. In bulk, there are one NN hop and two inequivalent 2NN hops: a 2NN hop perpendicular to the ferroelectric polarization direction, and a 2NN hop in a direction that forms an angle of ° with the polarization direction, see Figure 2(a). At ferroelectric domain walls, Fe sites are still equivalent within planes parallel to the wall, but differ between these planes. Hops can occur along the wall or with a component perpendicular to the wall, see Figure 2(b). The NEB calculations were considered converged when the energy barrier changed by maximally 1 meV in two consecutive iterations of the NEB structure optimization. The strain dependence of the barriers was also investigated. The strain reference (zero strain, ) corresponds to the optimized geometry of bulk or domain wall without excess electron. Tensile strain, , corresponds to the optimized geometry of bulk or domain wall with an excess electron.
@@ 2.4 Transition state theory
is the activation energy (hopping barrier) for the 2NN hop, is the reorganization energy. For symmetric hops, . The electronic coupling matrix elements , where the are electronic wave functions localized on the initial and final Fe site, respectively, were adopted from Ref. [40], where they were calculated for Fe(H2O) complexes as a function of the Fe-Fe distance from a model Hamiltonian that treated Fe with a pseudopotential and H2O as classical point charges. Rosso and coworkers obtain a coupling matrix element of 60 meV for second-nearest neighbors in hematite [41], where the intersite distance is smaller and electronic intersite coupling should be larger than in BiFeO3. Here, the matrix element for the second-nearest Fe-Fe distance calculated for bulk BiFeO3, 5.516 , was used throughout (-101.875 cm eV). Alternatively, the coupling matrix elements could in principle be determined as the deviation of the calculated barrier from the parabolic crossing point, see Figure 1(b). In this case one obtains a coupling matrix element of the order of meV, see Supp. Inf., with a large error bar. This result is not used here.
@@
where is the elementary charge, is the hopping distance, is the number of equivalent available final states for a given initial state, is the dimension, is Boltzmann’s constant, and is the temperature, here 300 K. For bulk, =3, =12 for 2NN hops, is the 2NN Fe-Fe distance, and the pseudocubic lattice constant (about 4 Å). For two-dimensional transport within a domain-wall plane, . Hops along the coordinate consist of two consecutive hops, such as or , with and . Hops along the coordinate can be single 2NN hops with and =2, such as , or double 2NN hops with and , such as or . The two-dimensional -type mobility and the diffusion constant in the domain wall are computed as
@@ 3 Results
3.1 Material properties of BiFeO3 obtained with different exchange-correlation functionals
Table 1 contains properties of BiFeO3 in the cubic and the rhombohedral phase obtained with different exchange-correlation functionals and from experiment. The calculations were performed using the primitive rhombohedral 10-atom unit cell of BiFeO3. A cutoff energy of 520 eV, a -point mesh of points, and energy convergence thresholds of eV and eV for electronic and atomic structure were employed, respectively, except for HSE06, where a -point mesh of points was used. The ionic polarization was computed as a weighted sum of atomic displacements,
Here, denotes atoms in the unit cell, and are actual positions in the ferroelectric (FE) and reference positions in the paraelectric (PE) phase, the charges are nominal ionic charges (Bi3+, Fe3+, O2-), the are weighting factors, and is the unit-cell volume.
| (eV) | (Å) | (Å) | (°) | (C/cm2) | |
| HSE06 | -1.15 | 5.48 | 5.63 | 59.15 | 73 |
| LSDA | -0.55 | 5.26 | 5.43 | 60.53 | 63 |
| LSDA+ | -0.90 | 5.44 | 5.52 | 59.82 | 61 |
| PBE | -0.73 | 5.54 | 5.59 | 59.44 | 52 |
| PBE+ | -0.84 | 5.56 | 5.71 | 59.05 | 68 |
| PBEsol | -1.00 | 5.46 | 5.54 | 59.93 | 66 |
| expt. | 5.64 | 59.42 | 72 |
Taking PBEsol or HSE06, which yield excellent structural agreement with experiment, as a benchmark, LSDA+ underestimates the ferroelectric energy gain by about 10–20%.
@@ 3.2 Excess electrons in bulk
| LSDA+ | HSE06 | |
| (eV) | -0.34 | -0.57 |
| (eV) | -0.71 | -1.2 |
The formation energy of small polarons is strongly negative (LSDA: 0.3 eV, HSE: 0.6 eV, see Table 2), hence the small polaron is the ground state and should be stable not only at low temperature, but also at room temperature and above. The electronic density of states (DOS) for a supercell (80 atoms) with an electron polaron is shown in Figure 3. The DOS obtained with LSDA+ and HSE06 agree semiquantitatively. The computed polaron level lies 0.7 eV below the conduction band minimum for LSDA+ and 1 eV below the conduction band minimum for HSE06, with an estimated uncertainty of 0.2 eV. Hence the polaron level is a deep level at room temperature.
The moderate strain dependence of polaron formation energy and electronic eigenvalue is depicted in Figure 4.
Hopping barriers
The energy barriers for NN hops and the two different 2NN hops are depicted in Figure 5(a) and (b) for different strains . Both NN and 2NN barriers are about 0.2 eV and slightly increase under tensile strain. Hence, the hopping barrier is considerably smaller than the energy distance of about 0.7 eV to the conduction band minimum. The 2NN barriers are not strongly direction dependent. The transitions from initial to final site are non-adiabatic (abrupt), as seen from the isosurface of the excess electron’s density as a function of the hopping coordinate (yellow orbitals in Figure 5(c) and (d)). The computed energy barriers are listed in Table 3.
| NN | 2NN, ° | 2NN, ° | ||||||
| LSDA+ | HSE06 | LSDA+ | HSE06 | LSDA+ | HSE06 | (cm2/Vs) | (cm2/s) | |
| % | 0.15 | 0.23 | 0.16 | 0.25 | 0.16 | 0.24 | – | – |
| 0 | 0.16 | 0.19 | 0.18 | 0.20 | 0.17 | 0.21 | – | – |
| 0.64% | 0.17 | 0.26 | 0.18 | 0.22 | 0.18 | 0.22 | ||
| 1.9% | 0.19 | – | 0.18 | – | 0.20 | – | ||
The barrier obtained with HSE06 is considerably larger, however the structure was not optimized with HSE06 but with LSDA+, therefore the barrier from HSE06 should be overestimated. In the following, the barrier obtained with LSDA+ is used. To obtain the electron mobility and the diffusion constant at room temperature, the extrapolated barrier at 1.9%, corresponding to the experimental lattice constant at room temperature [42], was used. With an energy barrier of 0.19 eV (the average of the two 2NN barriers) and a hopping distance of the 2NN Fe-Fe distance, the computed electron mobility in bulk BiFeO3 at room temperature is cm2/Vs, the computed diffusion constant is cm2/s. With an estimated uncertainty of 10% in the activation barrier, one obtains to cm2/Vs and to cm2/s.